OpenStudy (anonymous):

Find the difference of 9 radical 7 and radical 28. A.7 radical 7 B.11 radical 7 C.5 radical 7

OpenStudy (anonymous):

\[9\sqrt{7} - \sqrt{28}\] ?

OpenStudy (anonymous):

so C

OpenStudy (saifoo.khan):

\[7\sqrt{7}\]

OpenStudy (anonymous):

\[9\sqrt{7} and \sqrt{28}\]

OpenStudy (anonymous):

oh lol 9-2 = 7 so A

OpenStudy (saifoo.khan):

LOl.

OpenStudy (anonymous):

\[9\sqrt{7} - \sqrt{4*7} = 9\sqrt{7} - 2\sqrt{7} = 7\sqrt{7}\]